# LLM Agents That Act Like Us: Accurate Human Behavior Simulation with Real-World Data

### v4

@@ Abstract
Recent research shows that LLMs can simulate “believable” human behaviors to power LLM agents via prompt-only methods. In this work, we focus on evaluating and improving LLM’s objective “accuracy” rather than the subjective “believability” in the web action generation task, leveraging a large-scale, real-world dataset collected from online shopping human actions. We present the first comprehensive quantitative evaluation of state-of-the-art LLMs (e.g., DeepSeek-R1, Llama, and Claude) on the task of web action generation. Our results show that fine-tuning LLMs on real-world behavioral data substantially improves their ability to generate actions compared to prompt-only methods. Furthermore, incorporating synthesized reasoning traces into model training leads to additional performance gains, demonstrating the value of explicit rationale in behavior modeling. This work establishes a new benchmark for evaluating LLMs in behavior simulation and offers actionable insights into how real-world action data and reasoning augmentation can enhance the fidelity of LLM agents.
@@
LLM Agents That Act Like Us: Accurate Human Behavior Simulation with Real-World Data
Yuxuan Lu1,2, Jing Huang1, Yan Han1, Sisong Bei1, Yaochen Xie1, Dakuo Wang1,2, Zheshen Wang1, Qi He1 1Amazon.com, Inc., 2Northeastern University Correspondence: lu.yuxuan@northeastern.edu, d.wang@northeastern.edu
@@ 1 Introduction
Recent advances in large language models (LLMs) have enabled the simulation of “believable” human behavior across a range of applications, including web navigation actions Gur et al. (2023); Zhou et al. (2024), social interaction behaviors Park et al. (2023), interpersonal trust behaviors Xie et al. (2024), and user interface interactions Taeb et al. (2024). These developments have sparked growing interest in using LLMs to power agentic systems that emulate human behaviors (LLM Agent) Chen et al. (2025). Despite these successes, current evaluations Park et al. (2023) primarily emphasize subjective measures of “believability” (“how much people feel it is like a human”) instead of the objective “accuracy” (“how much it acts like a human”). The most relevant works that measure the objective model accuracy focus only on the final outcome of a task (e.g., purchasing the final product or not Yao et al. (2023), or ultimately trusting the partner or not Xie et al. (2024)), rather than rigorously evaluating whether the generated sequence of actions mirrors authentic human behavior or not. Consequently, the field currently lacks a robust and quantitative understanding for assessing LLMs at the process-centric, action-level simulation of human behaviors.
@@
To bridge this gap, our work provides the first systematic evaluation of SOTA LLMs’ accuracy in process-centric, action-level behavior simulation tasks. We leverage a large-scale, real-world dataset consisting of 31,865 user sessions of 3,526 users from an online shopping platform. Each session (Figure 1) comprises a series of timestamp-aligned <context, action> pairs, where the context reflects the webpage observed by the user (e.g., product views, filter states), and the action denotes user inputs such as clicks, searches, or session termination actions. In total, the dataset has 230,965 user actions, and the final outcomes of the sessions include 4,432 purchase actions and 27,433 session termination actions. This dataset enables us to rigorously evaluate how accurately various LLMs can generate human-like behaviors at the action level.
Beyond evaluation, our dataset uniquely positions us to fine-tune open-source LLMs to enhance their accuracy in behavior simulation tasks. While prior work has primarily relied on prompt-based approaches, we show that model adaptation through fine-tuning leads to significantly better accuracy in action generation and outcome prediction. Furthermore, drawing inspiration from reasoning-augmented modeling DeepSeek-AI et al. (2025), we hypothesize that exposing models to intermediate reasoning traces–even if synthetically generated–can enhance their ability to simulate human behavior. To test this hypothesis, we augment our dataset with synthesized reasonings from action traces using Claude 3.5 Sonnet and fine-tune models using this augmented data (<context, action, reasoning> triplets) to learn how to generate not only accurate actions but also the underlying reasoning. Our results show that this reasoning-augmented fine-tuning further boosts model performance, highlighting the importance of modeling not just what humans do, but also why they do it.
@@
We provide the first quantitative, process-centric evaluation of SOTA LLMs for simulating human web-action behaviors using real-world data from an online shopping context.
We empirically demonstrate that fine-tuning LLMs with synthesized reasoning traces significantly improves action-generation accuracy, highlighting the value of incorporating explicit reasoning to improve the realism and generalizability of behavior-simulating LLM agent applications.
@@ 2.1 Simulation of Human Behavior with LLM
The core function of the emerging LLM agent systems is their capability of generating human behaviors, in which a model takes a static user persona (e.g., preferences, demographics, or shopping habits), and the session data (e.g., a sequence of actions) as input to generate the next user action. Such systems have been extensively utilized and tested to simulate human behavior in a variety of scenarios. Park et al. (2023) simulated social behavior using generative agents in a virtual town, producing “believable” interactions. Xie et al. (2024) studied LLM agents in Trust Games to assess their ability to model human trust behavior. Park et al. (2024) used LLMs to simulate responses from 1,052 individuals in a social science survey. To simulate UI interaction, Lu et al. (2025) proposed UXAgent, enabling LLMs to operate within web environments for simulated usability testing. Collectively, these studies underscore the growing potential of LLM-driven simulations to model and simulate complex, interpretable human behaviors.
However, the evaluation of these works remains limited in scope. Some focus on the subjective believability of process-centric action traces. For instance, Lu et al. (2025) conducted qualitative interviews to assess participants’ perceptions of the realism of their UXAgent system. Similarly, Park et al. (2023) proposed an evaluation framework that identified emergent social behaviors among generative agents. On the other hand, works that pursue objective evaluation often do so in a single-shot, outcome-centric manner. Zhou et al. (2024) introduced WebArena, a controlled environment for benchmarking web agents based on task completion rates. ReAct Yao et al. (2023) measured success rates in simulation environments, overlooking the accuracy of step-by-step behavior. To date, no prior work has focused on objectively evaluating process-centric action traces—that is, assessing whether a model’s sequence of decisions faithfully aligns with human behavior at each step.
@@ 2.2 Reasoning in Human Behavior Simulation
However, the aforementioned works incorporate reasoning using prompt-only approaches for the action generation task. Whether reasoning can improve performance in fine-tuning settings remains an open question. More critically, there is currently no ready-to-use dataset designed specifically for human behavior simulation that includes both reasoning traces and corresponding actions. To address a similar cold-start problem in reinforcement learning, DeepSeek-AI et al. (2025) constructed a small-scale dataset of long reasoning traces by synthesizing examples through few-shot prompting, followed by reflection and verification. Inspired by their methodology, we apply a similar strategy to synthesize reasoning traces in our domain. This enables us to investigate whether integrating reasoning into fine-tuning can enhance a model’s ability to simulate human behavior more accurately.
@@ 3.1 Task Definition
At each time step $t$, the model is tasked with generating both the reasoning $r_{t}$ and the next action $a_{t}$. The model input includes the current context $c_{t}$ (what the user currently observes), a sequence of previous contexts (what the user has observed) $c_{1\dots t-1}$, a sequence of previous actions $a_{1\dots t-1}$ (what the user has done), and the corresponding reasoning trace $r_{1\dots t-1}$ (why the user had that action) within the same session. Formally, the model learns a function $f$ such that:
@@
Context
The Context, also referred to as the “observation space” of the web environment, represents the information available on a web page, including textual content, metadata, visual elements, and structural data. This context is designed to capture how users observe web pages, enabling an agent to process relevant information for navigation, retrieval, and interaction.
Following recent work Lu et al. (2025), we adopt a structured representation of the web page using a simplified HTML format, which preserves key structural elements while filtering out irrelevant details such as scripts and styling information. To facilitate precise action execution, each interactable element (such as links, buttons, and input fields) is assigned a unique hierarchical “name” that incorporates parent-child relationships within the webpage. These names serve as identifiers in the action space, ensuring that the agent can correctly reference and interact with specific elements. A description definition of our context design can be found in Appendix B.
@@
Previous research has explored various approaches to defining action spaces, including task-specific semantic actions such as “searching”, “adding items to a cart”, and “making purchases” Yao et al. (2022), as well as browser-level interactions like “typing” and “clicking” Lu et al. (2025).
@@ 3.2 Synthesized Reasoning Trace
The LLM is then prompted to generate a free-text reasoning explaining the user’s decision. This approach ensures that the reasoning traces are both coherent with the action and reflective of human reasoning. By bridging the gap between raw behavioral data and the reasoning processes, our method enhances the explainability of the model.
@@ 3.3 Model Architecture
To incorporate these enriched action traces, we build on existing pre-trained LLMs as our base models. The input to the model consists of two components: (1) a sequence of historical contexts (what the user observed), and (2) the corresponding actions and reasoning trace (what the user did and why). The model is trained to generate the next reasoning and action, enabling it to simulate human behavior with greater fidelity.
During the training stage, the model receives the full sequence of a user session—including context, reasoning, and action—as a single concatenated input. The training objective is to minimize the next-token prediction loss for the reasoning and action tokens, while the loss for the context tokens is masked out. Subsequently, in the evaluation stage, the model is provided with historical context and past reasoning traces and actions. The model first generates the reasoning for the next action; then, based on the generated reasoning, it generates the next action.
@@ 4.1 Dataset Construction
Our training dataset was constructed using data from one of the largest e-commerce platforms globally. The dataset contains 31,865 user sessions from 3,526 users in the online shopping scenario, comprising 230,965 user actions. The session’s final outcome includes 4,432 purchase actions and 27,433 session termination actions. We leveraged our data synthesis pipeline detailed in Section 3.2 to generate the synthesized reasoning for each action based on the context using Claude-3.5-Sonnet.
@@ 4.2 Evaluation and Metrics
We used a subset of the dataset that was not used during training as the test set, ensuring that no user sessions in the test set were seen by the model during fine-tuning. To create test cases, we took the second and all subsequent actions within each session—excluding the first action, since it lacks any preceding context. For each test case, the model was provided with the historical context, along with all previous actions and reasoning trace in the same session, and tasked with predicting both the next reasoning and the corresponding action.
This setup allows us to evaluate the model’s performance in a sequential reasoning setting, closely mirroring how real users behave during the session. This process resulted in a total of 932 test cases.
Baseline Method
@@
We evaluate model performance across two key dimensions: final outcome prediction and action generation.
@@
For action generation accuracy, we evaluate whether the model can accurately generate user actions during the process. We use an exact match criterion, where a generation is considered correct only if the action type, action target (e.g., search box or product link), and action attribute (if applicable, such as search keyword) exactly match the ground truth. To obtain the final score, we first compute the average per-session accuracy and then take the mean across all sessions to ensure a fair evaluation across different sessions and behaviors.
@@ 4.3 Experimental Setup
All fine-tuned models were trained using the same dataset and pipeline to ensure a fair comparison. Hyperparameters—including learning rate, batch size, and the number of training epochs—can be found in Appendix A.
@@ 4.4 Evaluation Results and Analysis
We then fine-tuned LLaMA 3.2, Qwen 2.5, and Mistral using our training dataset. The results demonstrate that fine-tuning LLMs with action traces and synthesized reasoning traces significantly enhances performance. Qwen 2.5-7B achieved 17.26% accuracy in action generation, surpassing DeepSeek-R1 by 5.4%. Similarly, LLaMA 3.2-3B reached an F1 score of 33.99% on the final outcome prediction task, further confirming the effectiveness of fine-tuning. These findings underscore that incorporating domain-specific fine-tuning with synthesized reasoning traces leads to substantial improvements in the accuracy of human behavior simulation.
We also visualize the distribution of user actions across different models compared to real human behavior in Figure 2. The fine-tuned models exhibit a distribution that more closely aligns with human action patterns, capturing a more natural balance between search, product clicks, and filter actions. In contrast, Claude 3.5 Sonnet displays a noticeably higher purchase rate while conducting only a single search per session, diverging from actual user behavior. This misalignment suggests that fine-tuning not only improves accuracy but also helps models better replicate realistic shopping behaviors rather than over-prioritizing purchase actions.

### v5

# Prompting is Not All You Need! Evaluating LLM Agent Simulation Methodologies with Real-World Online Customer Behavior Data

@@ Abstract
Recent research shows that LLMs can simulate “believable” human behaviors to power LLM agents via prompt-only methods. In this work, we focus on evaluating LLM’s objective “accuracy” rather than the subjective “believability” in simulating human behavior, leveraging a large-scale, real-world dataset collected from customers’ online shopping actions. We present the first comprehensive evaluation of state-of-the-art LLMs (e.g., DeepSeek-R1, Llama, and Claude) on the task of web shopping action generation. Our results show that out-of-the-box LLM-generated actions are often misaligned with actual human behavior, whereas fine-tuning LLMs on real-world behavioral data substantially improves their ability to generate accurate actions compared to prompt-only methods. Furthermore, incorporating synthesized reasonings into model training leads to additional performance gains, demonstrating the value of explicit rationale in behavior modeling. This work evaluates state-of-the-art LLMs in behavior simulation and provides actionable insights into how real-world action data can enhance the fidelity of LLM agents.
@@
Prompting is Not All You Need! Evaluating LLM Agent Simulation Methodologies with Real-World Online Customer Behavior Data
Yuxuan Lu1,2, Jing Huang1, Yan Han1, Bingsheng Yao2, Sisong Bei1, Jiri Gesi1, Yaochen Xie1, Zheshen (Jessie) Wang1, Qi He1, Dakuo Wang1,2, 1Amazon.com, Inc., 2Northeastern University Correspondence: lu.yuxuan@northeastern.edu, d.wang@northeastern.edu
@@ 1 Introduction
Recent advances in large language models (LLMs) have enabled the simulation of “believable” human behavior across a range of applications, including web navigation actions Gur et al. (2023); Zhou et al. (2024), social interaction behaviors Park et al. (2023), interpersonal trust behaviors Xie et al. (2024), and user interface interactions Taeb et al. (2024). These developments have sparked growing interest in using LLMs to power agentic systems that emulate human behaviors (LLM AgentChen et al. (2025). Despite these successes, current evaluations Park et al. (2023) primarily emphasize subjective measures of “believability” (“how much people feel it is like a human”) instead of the objective “accuracy” (“how much it acts like a human”). The most relevant works that measure the objective model accuracy focus only on the final outcome of a task (e.g., purchasing the final product or not Yao et al. (2023), or ultimately trusting the partner or not Xie et al. (2024)), rather than rigorously evaluating whether the generated sequence of actions mirrors authentic human behavior or not. Consequently, the field currently lacks a robust and quantitative understanding for assessing LLMs at the process-centric, action-level simulation of human behaviors.
@@
To bridge this gap, our work provides the first systematic evaluation of SOTA LLMs’ accuracy in process-centric, action-level behavior simulation tasks. We leverage a large-scale, real-world dataset consisting of 31,865 user sessions of 3,526 users from an online shopping platform. Each session (Figure 1) comprises a series of timestamp-aligned <context, action> pairs, where the context reflects the webpage observed by the user (e.g., product views, filter states), and the action denotes user inputs such as clicks, searches, or session termination actions. In total, the dataset has 230,965 user actions, and the final outcomes of the sessions include 4,432 purchase actions and 27,433 session termination actions. This dataset enables us to rigorously evaluate how accurately various LLMs can generate human-like behaviors at the action level. We evaluate different models on the next action generation task, benchmarking both the accuracy of generated actions throughout the session and the F1 score for final action generation, following protocols similar to existing work on modeling online shopping behavior.
Beyond evaluation, our dataset uniquely positions us to fine-tune LLMs to enhance their accuracy in behavior simulation tasks. While prior work has primarily relied on prompt-based approaches, we show that model adaptation through fine-tuning leads to significantly better accuracy in action generation and session outcome prediction. Furthermore, drawing inspiration from reasoning-augmented modeling DeepSeek-AI et al. (2025), we hypothesize that exposing models to intermediate reasoning traces–even if synthetically generated–can enhance their ability to simulate human behavior. To test this hypothesis, we augment our dataset with synthesized reasonings from action traces using Claude 3.5 Sonnet and fine-tune models using this augmented data (<context, action, reasoning> triplets) to learn how to generate not only accurate actions but also the underlying reasoning. Our results show that this reasoning-augmented fine-tuning further boosts model performance, highlighting the importance of modeling not just what humans do, but also why they do it.
@@
  • We provide the first quantitative and process-centric evaluation of LLMs to simulate human web action behaviors using real-world data from an online shopping context.
  • Through our evaluation, we empirically demonstrate that fine-tuning LLMs with synthesized reasoning significantly improves model’s performance, highlighting the value of incorporating explicit reasoning to improve the fidelity of LLM agent applications for role-playing or behavior simulation .
@@ 2.1 Simulation of Human Behavior with LLM
The core function of the emerging LLM agent systems is their capability of generating human behaviors, in which a model takes a static user persona (e.g., preferences, demographics, or shopping habits), and the session data (e.g., a sequence of actions) as input to generate the next user action. Such systems have been extensively utilized and tested to simulate human behavior in a variety of scenarios. Park et al. (2023) simulated social behavior using generative agents in a virtual town, producing “believable” interactions. Xie et al. (2024) studied LLM agents in Trust Games to assess their ability to model human trust behavior. Park et al. (2024) used LLMs to simulate responses from 1,052 individuals in a social science survey. To simulate UI interaction, Lu et al. (2025a) proposed UXAgent, enabling LLMs to operate within web environments for simulated usability testing. Collectively, these studies underscore the growing potential of LLM-driven simulations to model and simulate complex, interpretable human behaviors.
However, the evaluation of these works remains limited in scope. Some focus on the subjective believability of process-centric action traces. For instance, Lu et al. (2025a) conducted qualitative interviews to assess participants’ perceptions of the realism of their UXAgent system. Similarly, Park et al. (2023) proposed an evaluation framework that identified emergent social behaviors among generative agents. On the other hand, works that pursue objective evaluation often do so in a single-shot, outcome-centric manner. Zhou et al. (2024) introduced WebArena, a controlled environment for benchmarking web agents based on task completion rates. ReAct Yao et al. (2023) measured success rates in simulation environments, overlooking the accuracy of model mimicing step-by-step human behavior. To date, no prior work has focused on objectively evaluating model-generated action traces–that is, assessing whether a model’s sequence of decisions faithfully aligns with human behavior at each step.
@@ 2.2 Reasoning in Human Behavior Simulation
However, the aforementioned works incorporate reasoning using prompt-only approaches for the action generation task. Whether reasoning can improve performance in fine-tuning settings remains an open question. Although datasets for reasoning behind model actions exist (e.g., for conversational agents (Dongre et al., 2025)), there is currently no ready-to-use dataset specifically designed for human behavior simulation that includes both human actions and the corresponding reasoning behind those actions. To address a similar cold-start problem in reinforcement learning, DeepSeek-AI et al. (2025) constructed a small-scale dataset of long reasoning traces by synthesizing examples through few-shot prompting, followed by reflection and verification. Inspired by their methodology, we apply a similar strategy to synthesize reasoning traces in the online shopping domain. This enables us to investigate whether integrating reasoning into fine-tuning can enhance a model’s ability to accurately simulate human behavior.
@@ 3.1 Task Definition
At each time step $t$, the model is tasked with generating both the reasoning $r_{t}$ and the action $a_{t}$. The model input includes the current context $c_{t}$ (what the user currently observes), a sequence of previous contexts (what the user has observed) $c_{1\dots t-1}$, a sequence of previous actions $a_{1\dots t-1}$ (what the user has done), and the corresponding synthetic reasoning trace $r_{1\dots t-1}$ (why the user had that action) within the same session. Formally, the model learns a function $f$ such that:
@@
Observation Context
The Context (or “observation space”) of the web agent encompasses all available information on a webpage, including textual content, metadata, visual elements, and structural data. This context is designed to reflect how a human perceives and interprets a webpage, allowing the agent to process relevant features and perform tasks such as navigation, information retrieval, and interaction with page elements.
Previous research has explored various approaches to defining the context. Some methods rely on manually structured information parsing Yao et al. (2022), while others utilize raw HTML representations Gur et al. (2023) or accessibility tree embeddings Zhou et al. (2024). However, both approaches have notable limitations: manually structured information parsing requires significant human effort to develop parsing rules Yao et al. (2022), while raw HTML representations often contain extraneous information (e.g., CSS) that is not directly observable to human users, which may bias the LLM performance.
To ensure the adaptability to unseen websites, we define and implement a simplified HTML format as our context representation. This format removes non-relevant elements such as scripts, CSS, and purely visual components while preserving essential structural information. Using simplified HTML offers several advantages over custom formats or markdown-based representations: (1) important structural elements, such as lists and tables, remain intact, and (2) LLMs are already familiar with the HTML format, eliminating the need to redefine common elements like ‘button’ and ‘input’.
[…]
@@
Previous research has explored various approaches to defining action spaces, including task-specific semantic actions such as “searching”, “adding items to a cart”, and “making purchases” Yao et al. (2022), as well as browser-level interactions like typing and clicking Lu et al. (2025a).
@@ 3.2 Synthesized Reasoning Trace
The LLM is then prompted to generate a free-text reasoning explaining the user’s decision. This approach ensures that the synthetic reasoning traces are both coherent with the action and reflective of human reasoning. By bridging the gap between raw behavioral data and the reasoning processes, our method enhances the explainability of the model.
@@ 3.3 Model Architecture
To incorporate these enriched action traces, we build on existing pre-trained LLMs as our base models. The input to the model consists of two components: (1) a sequence of historical contexts (what the user observed), and (2) the corresponding actions and generated reasoning trace (what the user did and why). The model is trained to generate the next reasoning and action, enabling it to simulate human behavior with greater fidelity.
During the training stage, the model receives the full sequence of a user session—including context, synthetic reasoning, and action—as a single concatenated input. The training objective is to minimize the next-token prediction loss for the reasoning and action tokens, while the loss for the context tokens is masked out. Subsequently, in the evaluation stage, the model is provided with historical context and past reasoning traces and actions. The model first generates the reasoning for the next action; then, based on the generated reasoning, it generates the next action.
@@ 4.1 Dataset Construction
Our training dataset was constructed using data from one of the largest e-commerce platforms globally. The dataset contains 31,865 user sessions from 3,526 users in the online shopping scenario, comprising 230,965 user actions. The session’s final outcome includes 4,432 purchase actions and 27,433 session termination actions. We leveraged our data synthesis pipeline detailed in Section 3.2 to generate the synthetic reasoning for each action based on the context using Claude-3.5-Sonnet.
@@ 4.2 Evaluation and Metrics
We used a subset of the dataset that was not used during training as the test set, ensuring that no user sessions in the test set were seen by the model during fine-tuning. To create test cases, we took the second and all subsequent actions within each session, excluding the first action, since it lacks any preceding context. For each test case, the model was provided with the historical context, along with all previous actions and corresponding synthetic reasoning traces in the same session. The model is first tasked with predicting the next reasoning, and then, based on the reasoning it generated, it produces the corresponding action.
This setup allows us to evaluate the model’s performance in a sequential reasoning setting, closely mirroring how real users behave during the session. This process resulted in a total of 357 sessions with 932 test cases.
Prompt-Only Baseline
@@
We evaluate model performance across two key dimensions: Next Action Generation and Session Outcome Classification.
For Next Action Generation, we evaluate whether the model can accurately generate user actions during the process. We use an exact match accuracy, where a generation is considered correct only if the action type, action target (e.g., search box or product link), and action attribute (if applicable, such as search keyword) exactly match the ground truth. To avoid skewing the results toward longer sessions, we compute per-session accuracy first and then average across sessions, ensuring each session contributes equally to the final score.
@@ 4.3 Experimental Setup
All fine-tuned models were trained using the same dataset and pipeline to ensure a fair comparison.
We employed Fully Sharded Data Parallel Zhao et al. (2023) for efficient training. All sequences were padded or truncated to a context length of 40k tokens. We used a per-device batch size of 1, resulting in a global batch size of 64. The learning rate was set to 2e-5 with a cosine scheduler for adaptive learning rate adjustment. Models were trained for 1 epoch. For example, training Mistral-7B-v0.3 with a 40k token context window in our setup requires approximately 130GB of GPU memory per GPU, which is the largest model we can train.
Figure 3: Error Type Analysis of different models.
@@ 4.4 Evaluation Results and Analysis
We then fine-tuned LLaMA 3.2, Qwen 2.5, and Mistral using our training dataset. The results demonstrate that fine-tuning LLMs with action traces and synthesized reasoning traces significantly enhances performance. Qwen 2.5-7B achieved 17.26% accuracy in action generation, significantly surpassing DeepSeek-R1 by 5.4% ($p<10^{-10}$, McNemar’s test). Similarly, LLaMA 3.2-3B reached an F1 score of 33.99% on the final outcome prediction task, further confirming the effectiveness of fine-tuning. All fine-tuned models significantly outperformed their own ICL variants ($p<10^{-5}$, McNemar’s test). These findings underscore that incorporating domain-specific fine-tuning with synthesized reasoning traces leads to substantial improvements in the accuracy of human behavior simulation.
We also visualize the distribution of user actions across different models compared to real human behavior in Figure 2. The fine-tuned models exhibit a distribution that more closely aligns with human action patterns, capturing a more natural balance between search, product clicks, and filter actions. In contrast, Claude 3.5 Sonnet and DeepSeek-R1 displays a noticeably higher purchase rate while conducting only a single search per session, diverging from actual user behavior. This misalignment suggests that fine-tuning not only improves accuracy but also helps models better replicate realistic shopping behaviors rather than over-prioritizing purchase actions.